OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote at y = 0 and a vertical asymptote at, x = 2 and x = 1.

OpenStudy (anonymous):

Alright, when does an equation have a vertical asymptote?

OpenStudy (anonymous):

when the bottom equals zero

OpenStudy (anonymous):

correct. So if we want there to be a vertical asymptote at x=2 and x=1, we need an equation that has the bottom equal to zero when x=2 and x=1. Can you think of any?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

How about (x-2)(x-1)?

OpenStudy (anonymous):

oh is that the answer?

OpenStudy (anonymous):

That's the bottom part of the rational function, because if you plug in either x=2 or x=1, you get zero. Understand?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So now we have \[\frac{ a }{(x-2)(x-1) }\]

OpenStudy (anonymous):

When is there a horizontal asymptote?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

at zero

OpenStudy (anonymous):

For this problem, correct. There is a horizontal asymptote at y=0. Unlike vertical asymptotes, horizontal asymptotes can be crossed. A horizontal asymptote mainly indicates long term behavior a.k.a the limit of a function

OpenStudy (anonymous):

If the hor. asymptote= 0, then that means as x-> infinity and x-> -infinity, y-> 0

OpenStudy (anonymous):

ok so the answer is 0/x-2)(x-1

OpenStudy (anonymous):

No, not quite. If you have just zero in the numerator, then y always equals 0.

OpenStudy (anonymous):

You want to make the function have long term behavior of what I described above.

OpenStudy (anonymous):

so like 1-1

OpenStudy (anonymous):

To do that, you take the leading coefficients (the first numbers of the top and the bottom) and divide them by each other.

OpenStudy (anonymous):

so... right now we have \[\frac{ a }{ x^2}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

because x^2 is the first number if we expand (x-1)(x-2)

OpenStudy (anonymous):

what is a?

OpenStudy (anonymous):

a is the top part of the function that we are still trying to find.

OpenStudy (anonymous):

yeah what would it be how do you find it

OpenStudy (anonymous):

\[\frac{ a }{ x^2}=0\]

OpenStudy (anonymous):

x+1)(x+2?

OpenStudy (anonymous):

a=x+1)(x+2?

OpenStudy (anonymous):

That's an idea, but if we had that in the top then the leading coefficient becomes x^2

OpenStudy (anonymous):

Then we would have \[\frac{ x^2 }{ x^2 } \]

OpenStudy (anonymous):

ok so then what is it i dont get it

OpenStudy (anonymous):

It's pretty simple I'm just having a hard time explaining it. The top can just be x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

If we just have x, then we get \[\frac{ x }{ x^2}\]

OpenStudy (anonymous):

which as x gets really big equals 0

OpenStudy (anonymous):

so final answer is \[\frac{ x }{ (x-2)(x-1)}\]